Write down the parametric coordinates of any point on each of the following parabolas.
(a) y^2=8x
b) y^2 =24x
c) y^2 =-16x
d)x^2 =6y
1
Expert's answer
2021-07-25T10:26:24-0400

(a) The parametric coordinates of any point on parabola "y^2 = 8x" :


"\begin{cases}
 x = \dfrac {t^2} 8 \\
 y = t 
\end{cases} \"


"-\infin < t < \infin"


(b) The parametric coordinates of any point on parabola "y^2 = 24x":

"\begin{cases}
 x = \dfrac {t^2} {24} \\
 y = t 
\end{cases} \"


"-\infin < t < \infin"


(c) The parametric coordinates of any point on parabola "y^2 = - 16 x" :

"\begin{cases}
 x = -\dfrac {t^2} {16} \\
 y = t 
\end{cases} \"


"-\infin < t < \infin"


(d) The parametric coordinates of any point on parabola "x^2 = 6y" :


"\begin{cases}
 x = t \\
 y = \dfrac {t^2} 6 
\end{cases} \"


"-\infin < t < \infin"